OpenStudy (anonymous):

What is the value of "x" ... 1+3(1+3(1+3(1+x))) = 13

OpenStudy (anonymous):

no

OpenStudy (anonymous):

x=-1?

OpenStudy (mathmate):

Yes, for the initial question.

OpenStudy (anonymous):

3(1+3(1+3(1+x))) = 12 (1+3(1+3(1+x)) = 4 3(1+3(1+x)) = 3 (1+3(1+x)) = 1 3(1+x)) = 0 1+ x = 0 x = -1